Problem 2: How much charge is on each capacitor? E 10V, C1 3nF, C2 6nF, and C3 9nF
Answer #1

here we observed C1 and C2 are parallel connection

Cp=3+6=9nF

Cp and C3 are series connection

the equivalent capacitance Ceq=9*9/9+9=4.5 nF

step;2

now we find the total charge

the total charge Q=4.5*10^-9*10=4.5*10^-8 C

step;3

now we find the charge on each capacitor

charge on capacitor C3=Q3=4.5*10^-8 C

charge on capacitor C1=>Q1=3*4.5*10^-8/3+6=1.5*10^-8 C

charge on capacitor C2=>Q2=6*4.5*10^-8/3+6=3*10^-8 C